OpenStudy (anonymous):

please help! x^4-x^3-2x-4=0 how do i solve the equation by first finding the rational roots?

OpenStudy (anonymous):

possible roots are +-1, +-2, +-4

OpenStudy (anonymous):

then i found the roots -1 and 2 by synthetic division, how do i find the third/fourth root?

OpenStudy (anonymous):

factor, you will have a quadratic left, use the quadratic formula

OpenStudy (anonymous):

i have two quadratics left from -1 and 2. which one do i use?

OpenStudy (anonymous):

you know what i means by factor? you said -1 is a root, so \[x^4-x^3-2x-4=(x+1)(\text{some cubic})\]

OpenStudy (anonymous):

you can find the cubic by synthetic division

OpenStudy (anonymous):

then from the cubic expression factor out the other root \(x-2\) so you will have \[x^4-x^3-2x-4=(x+1)(x-2)(\text{some quadratic})\]

OpenStudy (anonymous):

perhaps you can factor the quadratic, but you can always set that equal to zero and solve via the quadratic formula

OpenStudy (anonymous):

I'll try solving it

OpenStudy (anonymous):

It works. Thank you! So this is the method for solving any question like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you should have got \(x^2+2\) for the quadratic

OpenStudy (anonymous):

my final answer is {-1,2, +-sqrt2i}